OpenStudy (anonymous):

Is it true that every even number n ≥ 12 can be written as the sum of two different primes plus a multiple of 4 (i. e. for all n ≥ 12, n = p + q + 4k, where p and q are primes, p does not equal q, and k ∈ N ∪ {0})? If so, prove it, if not, provide a counterexample.

OpenStudy (anonymous):

yes its true

OpenStudy (anonymous):

even numbers can be written as 4m or 4m+2 right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4m=3+5+4(m-2)

OpenStudy (anonymous):

can u find a similar thing for 4m+2 ???

OpenStudy (anonymous):

I dont understand why you have 4(m-2) what is m-2? Can you please explain

OpenStudy (anonymous):

k is a arbitrary number right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so when i want to write 4m in the form p+q+4k i choose k=m-2

OpenStudy (anonymous):

let suppose u have n=16=4*4 so m=4 i choose k=m-2=2 so 16=3+5+4*2

OpenStudy (anonymous):

oh i see thank you

OpenStudy (anonymous):

what about the other one?